Calculate the mole fraction of the solute in the following solutions.


a) 100g C2H6O in 100g H2O


b) 30% HCI solution by mass


1
Expert's answer
2022-03-13T15:17:03-0400


a) moles C2H6O = 100g / 46 = 2.174

moles H2O = 100g / 18 = 5.56

total = 2.174 + 5.56 = 7.734

% = 2.174 / 7.734 x 100% = 28%

b)moles HCl = 30g / 36.5 = 0.82

moles H2O = 70g / 18 = 3.88

total = 0.82 + 3.88 = 4.7

% = 0.82 / 4.7 = 0.174 x 100% = 17.4%